A local paper claims the proportion of college students who own a car is 38%. In a sample of 275 college students, what is the probability that more than 116 will own a car?


1
Expert's answer
2021-05-11T08:13:03-0400

"p = 0.38 \\

q = 0.62 \\

n = 275"

Using binomial distribution:

Mean:

"\mu = n \times p = 275 \times 0.38 = 104.5"

Standard deviation:

"\sigma = \sqrt{n \times p \times q} = \sqrt{275 \times 0.38 \times 0.62} = 8.05 \\

P(X>116) = 1 -P(X<116) \\

= 1 -P(Z< \frac{116-104.5}{8.05}) \\

= 1 -P(Z<1.428) \\

= 1 -0.9233 \\

= 0.0767"